OpenStudy (anonymous):

arccosx=arctanx prove sin(arccosx)=x^2

OpenStudy (anonymous):

m ans is comin x/sqrt1+x^2

OpenStudy (anonymous):

thts the condition for the second statement

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

theres no real soln for the first statemnt

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

there is actually

OpenStudy (anonymous):

see arccosx = arcsin(root(1-x^2))

OpenStudy (anonymous):

always....but in ur case its equal to x^2....so root(1-x^2) = x^2

OpenStudy (anonymous):

thts the first condition

OpenStudy (anonymous):

hain na?

OpenStudy (anonymous):

den shal v conlude quest is wrng

OpenStudy (anonymous):

if u open the first statement up u get....x^4 + x^2 - 1=0 which is equivalent to x^4 = 1 - x^2

OpenStudy (anonymous):

so proved....

OpenStudy (anonymous):

m nt getin :(

OpenStudy (anonymous):

dekho yaar..suno dhyan se

OpenStudy (anonymous):

kaise sunu

OpenStudy (anonymous):

\[\sin^{-1}\sqrt{1-x ^{2}} = \cos^{-1}x\] and the second statement says that \[\cos^{-1} x = \sin^{-1}x ^{2}\] so equating the two statements u can say that \[x ^{2}= \sqrt{1-x^{2}}\]

OpenStudy (anonymous):

him...i think the second equation to prove...how can u start with this?

OpenStudy (anonymous):

No looking at ur first statement \[\cos^{-1} x=\tan^{-1} x\] which means \[x =1 /\sqrt{1 + x^{2}}\] or \[x^{4} + x^{2} = 1\]

OpenStudy (anonymous):

this the second equation says

OpenStudy (anonymous):

i think smthin is wrng bcz its nt posible

OpenStudy (anonymous):

ab kya hua yaar?

OpenStudy (anonymous):

no no him is right...

OpenStudy (anonymous):

jai dipankar baba ki...lol:)

OpenStudy (anonymous):

u conti

OpenStudy (anonymous):

Set \(\cos^{-1}x=\tan^{-1}x=\theta\), then draw a triangle that satisfies the given conditions. A triangle is attached. From the triangle you can easily see that \(\sin(\cos^{-1}x)=\sin\theta=x^2\).

OpenStudy (anonymous):

so let me to do a few steps.. him got x^4=1-x^2 from first step... sin (arccosx)=sin (arcsinsqrt(1-x^2))=sin(arcsinsqrt(x^4))=sin(arcsin x^2)=x^2 this proves ur question...

OpenStudy (anonymous):

yes..now u got it?

OpenStudy (anonymous):

samajh ayi baat?

OpenStudy (anonymous):

ya agai

OpenStudy (anonymous):

Are you guys writing in Hindi? :D

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

nice to be homolingual in a big large world:)

OpenStudy (anonymous):

Anwar..dou know hindi?

OpenStudy (anonymous):

NO :(

OpenStudy (anonymous):

so ui know the term hindi...

OpenStudy (anonymous):

Yeah of course.

OpenStudy (anonymous):

Anwar come in chat box..